OpenStudy (anonymous):

solve, round to nearest hundreth if necessary 11+ 4x/-5 = 2/3

OpenStudy (anonymous):

11 + (4x/-5) = (2/3) (4x/-5) = (2/3) -11 (4x /-5) = (-31/3) 4x = (-31/3) * -5 4x = (155/3) x = (155/3) / 4 x = 155/12

OpenStudy (anonymous):

yup that's right